OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 3) and (2, 7)? x - y = -1 x - y = -2 x - y = - 5 x - y = - 6

OpenStudy (abb0t):

Start by finding the slope: \[slope = m = \frac{ y_2-y_1 }{ x_2-x_1 }\] then pick ONE of the two points and plug it in for y sub zero: \[(y-y_0)=m(x-x_0)\]

OpenStudy (anonymous):

x - y = -2

OpenStudy (anonymous):

is that it

OpenStudy (abb0t):

Not quite.

OpenStudy (abb0t):

What's your slope (m). Start with that.

OpenStudy (anonymous):

x - y = - 6

OpenStudy (abb0t):

No. Are you just guessing now or actually working on it? The answer is quite clear when you actually do it.

OpenStudy (anonymous):

x - y = -1

OpenStudy (abb0t):

Show me your steps.

OpenStudy (anonymous):

thats the answer i did it

OpenStudy (abb0t):

No. Because if you did it, you would see that it's not the correct answer.

OpenStudy (abb0t):

Also that fact that you guessed every possible answer EXCEPT the correct one shows that you didn't try it. \[slope = m = \frac{ 7-3 }{ 2-(-2) }= \frac{ 4 }{ 4 } = 1\] \[y-7=x-2 => x-y=-5\]

OpenStudy (aravindg):

@abbot you have him the answer away unknowingly !!

OpenStudy (aravindg):

@abb0t

OpenStudy (abb0t):

I gave up cuz i was getting frustrated. Lol @AravindG

OpenStudy (aravindg):

@abb0t a better idea would have been to report this user. I just went through the questions he asked and I was like "These all were happening in Openstudy and no mods noticed o.O "

OpenStudy (abb0t):

lol. my bad. thnx for the tip :) @AravindG